OpenStudy (jaylelile):

Help me please! I don't know how to do this.. write the equation of the line (in slope intercept form) that is parallel to the given line but which passes through the point (-3,4) Also write the equation of the line line (in slop intercept form) that is perpendicular to the given line but which passes through (-3, 4) the "given" line it's talking about is y=-2x+1 or (-4,7) and (1,-3)

OpenStudy (baru):

can you identify "slope " of the given line?

OpenStudy (baru):

this is the info you will need to solve this: parallel lines have the same slope; if two lines are perpendicular, the product of their slopes is -1

OpenStudy (jaylelile):

so the parallel line has the same slope as my original line? this being -2? so how do I write the equations in slope intercept form?

OpenStudy (baru):

Y=MX + c is the slope intercept form,

OpenStudy (baru):

M is slope

OpenStudy (jaylelile):

so for my parallel line it would be y=-2+1 and my perpendicular line would be y=-1+1?

OpenStudy (baru):

No, where did the x go?

OpenStudy (baru):

Do the parallel lines first

OpenStudy (jaylelile):

oops so parellel would be y=-2x+2 and perpendicular would be y=-1x+1?

OpenStudy (baru):

For the parallel line Y=-2x + c U have to substitute the point -3,4 to find c

OpenStudy (jaylelile):

I got y=-2x+1.5

OpenStudy (baru):

No Substitute x=-3 and y=4 And find c

OpenStudy (jaylelile):

I got y=-2x+2....

OpenStudy (baru):

Yea

OpenStudy (jaylelile):

so that's my parallel line. and my perpendicular line is y=-1x+1??

OpenStudy (baru):

No Product of slopes of perpendicular lines =-1

OpenStudy (jaylelile):

i'm lost...

OpenStudy (baru):

What is the slope of the given line?

OpenStudy (jaylelile):

-2

OpenStudy (baru):

Let slope of perpendicular line be 'm' -2*m=?

OpenStudy (jaylelile):

4....

OpenStudy (baru):

When the slope of perpendicular lines are multiplied, we get -1

OpenStudy (baru):

-2*m=?

OpenStudy (jaylelile):

I'm confused...

OpenStudy (baru):

What's confusing?

OpenStudy (jaylelile):

I just don't understand what my equation is for my perpendicular line,.

OpenStudy (baru):

To find that, u first need to find slope of the perpendicular line

OpenStudy (baru):

Thats what I'm asking u to find

OpenStudy (jaylelile):

I thought the slope was -1?

OpenStudy (baru):

How is the slope-1

OpenStudy (jaylelile):

I thought you said it was... how do I find the slope of that line?

OpenStudy (baru):

If two lines are perpendicular, the product of their slopes is -1

OpenStudy (jaylelile):

okay......?

OpenStudy (baru):

This means, slope of given line multiplied by slope of the line u want to find gives -1

OpenStudy (jaylelile):

I don't understand what to solve here...

OpenStudy (baru):

U have to find "slope" of the required line, U can find that by using the property: slope of required line multiplied by slope of given line =-1

OpenStudy (baru):

Slope of given line: -2 Slope of required line: m -2*m=-1

OpenStudy (jaylelile):

0.5???

OpenStudy (baru):

yes

OpenStudy (jaylelile):

so what do I do now?

OpenStudy (baru):

y=0.5x + c find c by substituting the point (-3,4)

OpenStudy (jaylelile):

y=0.5x+0.375?

OpenStudy (baru):

no

OpenStudy (jaylelile):

what am I doing wrong?

OpenStudy (baru):

substitute x=-3 y=4 can you do that and re-write the equation?

OpenStudy (jaylelile):

\[4=0.5(-3)+c\] 0.5 multiplied by -3 \[4=1.5+c\] Both sides divided by 4 \[-0.375c\]

OpenStudy (jaylelile):

I got 5.5

OpenStudy (jaylelile):

thanks for all your help!

OpenStudy (baru):

yes, its 5.5